OpenStudy (anonymous):

The extremes in the proportion 3/4=15/20 are 3 and 4 3 and 20 15 and 4 15 and 20

geerky42 (geerky42):

In the proportion \(\dfrac{a}{b} = \dfrac{c}{d}\), a and the d are called the extremes and the b and c are called the means

OpenStudy (anonymous):

so 3/4?

geerky42 (geerky42):

a is 3 and d is 20

geerky42 (geerky42):

\(\dfrac{3}{4} = \dfrac{15}{20}\)

OpenStudy (anonymous):

thanks